OpenStudy (anonymous):

What is 9/15 written as a percent? A. 0.6% B. 6% C. 60% D. 600%

OpenStudy (anonymous):

C.

OpenStudy (dejmccray):

what is 9 divided by 15?

OpenStudy (anonymous):

Nevermind, just divide 9 / 15 = 0.60. Which is the same as 60/100 which is the same thing as 60%.

OpenStudy (anonymous):

0.6

OpenStudy (anonymous):

C

OpenStudy (anonymous):

60%

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

no its C

OpenStudy (anonymous):

Help on the next?

OpenStudy (anonymous):

To the nearest tenth of a percent, what is 5/7 written as a percent? A. 714.3% B. 71.4% C. 7.1% D. 0.7%

OpenStudy (anonymous):

no you divide 9 divided 15 it 0.6

OpenStudy (anonymous):

b

OpenStudy (anonymous):

b

OpenStudy (anonymous):

B

OpenStudy (anonymous):

What is 8.75% written as a fraction in lowest terms? A. 8 3/4 B. 7/80 C. 8/25 D. 17/ 50

OpenStudy (anonymous):

5 / 7 = 0.714285714 = 71.4/100 = 71.4%

OpenStudy (anonymous):

Here is what it is for the first one \[9/15 = .6\] \[.6 \times 100 = 60%\]

OpenStudy (anonymous):

5 / 7 = 0.714285714 = 71.4/100 = 71.4%

OpenStudy (anonymous):

8 3/4 = 8.75

OpenStudy (anonymous):

So A for the last one?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

For the last one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

B? You sure?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

No, it's A.

OpenStudy (anonymous):

7 / 80 = 0.0875 = 0.87/100 = 87%

OpenStudy (anonymous):

dude you helped me with the last problem and got it WRONG

OpenStudy (anonymous):

8 and 3/4 = 8.75

OpenStudy (anonymous):

SOMEONE ELSE TELL ME WHAT THE ANSWER TO THIS ISS. What is 8.75% written as a fraction in lowest terms A. 8 3/4 b.7/80 c. 8/25 d.17/50

OpenStudy (anonymous):

7 / 80 = 87.5%

OpenStudy (anonymous):

7/80ths

OpenStudy (anonymous):

19x3 + (14x + 4x3) =

OpenStudy (anonymous):

So B?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

A cross country team is made up of 65% male runners. What fraction of the team is male? A.1/65 B.2/3 C.13/20 D.7/10

OpenStudy (anonymous):

D 7/10

OpenStudy (anonymous):

Any other answers?

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

A bank paid 2.3% interest on a money market account. What is 2.3% written as a decimal? A. 0.023 B. 0.23 C. 2.3 D. 23.0

OpenStudy (anonymous):

Last one for now.

OpenStudy (anonymous):

.-.

OpenStudy (anonymous):

A

OpenStudy (anonymous):

0.023

OpenStudy (anonymous):

Thank you I passed

OpenStudy (anonymous):

Good Job:)

OpenStudy (anonymous):

ANother 5 questions here we go.

OpenStudy (anonymous):

Your from k12?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

what grade?

OpenStudy (anonymous):

just curious

OpenStudy (anonymous):

You spin the spinner below once. What is the probability that the arrow will land on 3? A.1/8 b.1/4 c.3/8 d.8

OpenStudy (anonymous):

7th grade

OpenStudy (anonymous):

cool im 6th

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

A

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

you count how many spaces there are

OpenStudy (anonymous):

and there is only one 3

OpenStudy (anonymous):

qand 8 spaces total

OpenStudy (anonymous):

*and

OpenStudy (anonymous):

The arrow on the spinner is to be spun once. What is the probability that the arrow will stop on a number that is greater than 5? A.2/5 B.4/9 C.1/2 D.5/9

OpenStudy (anonymous):

A 2/5

OpenStudy (anonymous):

A bag contains 3 red marbles, 5 yellow marbles, and 4 blue marbles. One marble is chosen at random. What is the probability that the chosen marble will be blue? A.1/12 B.1/4 C.1/3 D.3/4

OpenStudy (anonymous):

i think its A

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

A box contains only sliver tokens and gold tokens. Clarisse reached into the bag without looking, got 60 tokens, and found that 48 of the tokens were silver. Clarisse is going to reach in the box again and get one more token. What is the experimental probability that the token will be silver? A.1/48 B.2/5 C.1/2 D.4/5

OpenStudy (anonymous):

i beilieve it's D

OpenStudy (anonymous):

it is d 4/5

OpenStudy (anonymous):

Thanks ;D

OpenStudy (anonymous):

hat is the distance between the points (4, 7) and (4, –5)? A. 2 units B. 12 units C. 0 units D. 4 units

OpenStudy (vianne):

I just want to clarify for the future people that.. A bag contains 3 red marbles, 5 yellow marbles, and 4 blue marbles. One marble is chosen at random. What is the probability that the chosen marble will be blue? A.1/12 <-- Is wrong.. B.1/4 C.1/3 D.3/4 It's 4/12, simplified is 1/3 :)